A girl of mass 25kg running with a speed of 1.5m/s jumps on a skateboard  of mass 2kg kept on smooth surface.  What will be their combined velocity? 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Apply consevation of momentum
m1v1+m2v2 = (m1+m2)v
25×1.5+2×0 = (25+2)v
v= 1.388
 
Regards
Arun (askIITians forum expert)